OpenStudy (anonymous):

What is the radius of a sphere with a volume of 5.66cm^2? V= 4/3 * pi * r^3

OpenStudy (anonymous):

5.66^2=4/3*pi*r^3

jimthompson5910 (jim_thompson5910):

Did you mean 5.66 cm^3 and not cm^2? If so, then V= 4/3 * pi * r^3 5.66 = 4/3 * pi * r^3 5.66*3 = 4 * pi * r^3 16.98 = 4 * pi * r^3 16.98/4 = pi * r^3 4.245 = pi * r^3 pi * r^3 = 4.245 r^3 = 4.245/pi r^3 = 4.245/3.14 I'll let you finish

OpenStudy (anonymous):

if you get pi=3.14 answer is 1.1057306519563767583656432691051

OpenStudy (anonymous):

but with real Pi answer is: 1.1055437675995148584792024538633

OpenStudy (anonymous):

If 4.245/3.14= 1.35191, how do you find the cube route of it?

jimthompson5910 (jim_thompson5910):

cube root of 1.35191 = (1.35191)^(1/3)

OpenStudy (anonymous):

And that's the radius?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can use 0.33 for 1/3 in (1.35191)^(1/3) (1.35191)^(0.33)

OpenStudy (anonymous):

because 1/3=0.33333333

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

your welcome